A silver teaspoon at temperature 7⁰C is placed in a cup of hot coffee at temperature 93⁰C. If the teaspoon has a mass of 32 g and the hot coffee has 90g of mass, determine the final temperature of the silver teaspoon? Ccoffee = 1 cal/g.K Csilver = 0.0558 cal/g.K


1
Expert's answer
2022-01-18T10:20:42-0500

Heat loss by Coffee = - Heat gain by silver teaspoon.

From "Q= mc\theta"

"90\u00d71\u00d7(366-T_2)=-32\u00d70.0558\u00d7(280-T_2)"


"=32940-90T_2=-499.97+1.79T_2"

"T_2=364.31K =91.31\u00b0C"